# Recurrent Drafter for Fast Speculative Decoding in Large Language Models

### v2

@@ 3.1 Training
We used pre-existing Medusa checkpoints from the authors’ repository, which use Vicuna 7B base model (Touvron et al., 2023; Chiang et al., 2023), to assess training efficacy. For our method, ReDrafter, we adopted an identical training scheme, using a two-epoch training of the drafting head based on ShareGPT (2023). We subsequently performed evaluations using Alpaca Eval (Dubois et al., 2023).
@@
We analyzed the relationship between training accuracy and inference speed across 30 drafters, divided into two categories: drafters without recurrent neural networks (RNN) and with RNN integration. For each category, we adjust the number of ResNet layers and the learning rate, yielding 15 distinct configurations per category. For the non-RNN drafters, we adopted a simplified model setting where $U=I$, $W=I$, $b=0$, and applying an identity function $f(x)=x$. During interence, we use greedy decoding and set beam width equals to 15 for a fair comparison. Results are shown in Figure 4. For the drafters without RNN, we observed a high Pearson correlation coefficient 0.907, between validation accuracy at $t+5$ and inference speed within this group. Drafters equipped with RNN showed a Pearson correlation coefficient of 0.798.
The difference between the two groups is visually represented by the gap between blue and yellow dots in our scatter plots. This gap underscores the trade-off associated with employing a more complex modeling approach (with RNN). Although the use of RNN introduces additional computational complexity, it also enables the models to potentially achieve a higher ceiling in terms of both accuracy and efficiency during inference.
@@ 3.3 Analysis
Figure 6: (a) Speed (tokens/second) and throughput (batch size * tokens/second) for AR and ReDrafter with various beams. (b) Speed and throughput for ReDrafter w, w/o tree attentaion. RD stands for ReDrafer.
@@ 4 Conclusions
Acknowledgments.
We would like to thank Barry Theobald, Frank Chu, Liangliang Cao, Ruoming Pang, Sam Wiseman, Xiujun Li and Zhiyun Lu for their comments and suggestions on the manuscript. We thank Tianle Cai for sharing the details of Medusa with us.

### v3

@@ 3.1 Training
We used one existing Medusa checkpoint from the authors’ repository, Vicuna 7B base model (Touvron et al., 2023; Chiang et al., 2023), to assess training efficacy. For our method, we used a similar two-epoch training of the drafting head based on ShareGPT (2023). We subsequently performed evaluations using Alpaca Eval (Dubois et al., 2023).
@@
We analyzed the relationship between training accuracy and inference speed across 30 drafters, divided into two categories: drafters without RNN and with RNN. For each category, we adjust the number of ResNet layers and the learning rate, yielding 15 distinct configurations per category. For the non-RNN drafters, we adopted a simplified model setting where $U=I$, $W=I$, $b=0$, and applying an identity function $f(x)=x$. During inference, we use greedy decoding and set beam width equals to $15$ for all experiments. Results are shown in Figure 4. For the drafters without RNN, we observed a high Pearson correlation coefficient 0.907, between validation accuracy at $t+5$ and inference speed within this group. Drafters equipped with RNN showed a Pearson correlation coefficient of 0.798. The difference between the two groups is visually represented by the gap between blue and yellow dots in our scatter plots. This gap underscores the trade-off associated with employing a more complex modeling approach (with RNN). Although the use of RNN introduces additional computational complexity, it also enables the models to potentially achieve a higher ceiling in terms of both accuracy and efficiency during inference.
@@ 3.3 Analysis
Figure 6: (a) Speed in tokens per second and throughput (batch size * tokens/second) for auto-regressive and ReDrafter with various beams. (b) Speed and throughput for ReDrafter with and without tree attention. AR stands for auto-regressive and RD for ReDrafer.